I need help to find out if this equation is identity or contradiction
3(x+2)-2=-(5+x)+x

Answer by 1998Baby(63)

3(x+2)-2=-(5+x)+x
first distribute the numbers outside of the parenthesis.
3x+6-2=-5-x+x
then you can combine whatever you can on their own sides
3x+4=-5
then subtract 4 to -5
3x=-9
divide 3 to -9
x=-3
____
it is identity